OpenStudy (anonymous):

A triangle has a base of 9 feet and a height of 5 feet. What is its area

OpenStudy (anonymous):

22.5

OpenStudy (amistre64):

the area of a triangle is half that of a square 9*5 equals the area of a square; so divide it in half

OpenStudy (anonymous):

Lets not give answers, shall we, just give ways to solve, hints tips and formulae!

OpenStudy (anonymous):

A=1/2 bh a= 1/2* 9*5

OpenStudy (anonymous):

@amistre64 square with sides of different measures?????

OpenStudy (anonymous):

half bh =to area 22.5

OpenStudy (amistre64):

lol ... you oughta see the squares I draw :)

OpenStudy (anonymous):

maybe u wanted to say half of the rectangle.....

OpenStudy (anonymous):

if a square has different measure then it is not called a square

OpenStudy (amistre64):

good thing I didnt say it was a hexagon ... whew! then noone would know what i was saying ;)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

it is not half of a rectantangle coz height is different from length or breadth

OpenStudy (amistre64):

all triangles that can be formed from a rectangle with base 9 and height 5 area precisely half the area of the rectangle. you simply slide the top part of the triangle across the top part of the rectangle and leave the base as is

OpenStudy (anonymous):

Only 2 triangles can be drawn, draw a diagonal and it ll disect the rectangle into its two triangles. Thus area the half

OpenStudy (amistre64):

well; at most 3 trianlges can be formed; 1 is half the area; the other 2 combine to make the other half.

OpenStudy (anonymous):

you can similarly divide the first half into its two constituents in which case it ll give 4 and you can go on until infinity :D :D

OpenStudy (amistre64):

can something have an infinite surface area but a finite volume?

OpenStudy (anonymous):

I don't think so, cause volume is an extension of the surface in the third dimension, so anything multiplied by infinity gives infinity.

OpenStudy (anonymous):

Any triangle can be converted into a parallelogram by drawing appropriate lines... and one of the sides of the triangle will become the diagonal of the resulting ||gm so the ||gm will consist of two triangles of equal area so area of the triangle will be half the area of ||gm area of a ||gm = base*height base*height so area of triangle = ------------ !!!!! 2

OpenStudy (amistre64):

gabriels horn; has infinite surface area with a volume of pi

OpenStudy (anonymous):

I am not aware of that amistre64, i ll study about it :)

OpenStudy (anonymous):

ok, i just read, but it can never have infinite surface area, and so the volume will always be less than pi, so at certain point we make an approximate.

OpenStudy (amistre64):

it can by definition: \[\int_{1}^{\infty}\pi(1/x)^2dx\implies\pi\text{ as a volume of revolution}\] whereas the infinite length of the surface area created by the revolution is infinite :)

OpenStudy (amistre64):

one might argue that pi itself is an infinite number; never ends. But we know that pi is between 3 and 4 units. such that the horn can be overfilled with paint; but never covered by paint ...

OpenStudy (anonymous):

Thats a very interesting way to look at it amistre64, using the paint example, no wonder you are in the group sensei :D, But i believe we are spoiling this question by discussing about something totally unrelated never the less very interesting :D